ABC Inc. recently issued $1,000 par bonds at a 2.1% coupon rate. If the bonds have 20 years to maturity and a YTM of 16.94%, what is the current price of the bond? Assume semi-annual compounding.Note: Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

As Semi-annual compounding,

Coupon Rate = 2.1%/2 = 1.05%

YTM = 16.94%/2 = 8.47%

Years = 20*2 = 40

Bond Price = Interest * PVIFA(YTM, Years) + Par Value * PVIF(YTM, Years)

= (1000*1.05%) * PVIFA(8.47%,40) + 1000 * PVIF(8.47%,40)

= 10.50 * 11.34957 + 1000 * 0.03869

= 119.1705 + 38.69

Bond Price = $ 157.86